OpenStudy (anonymous):

i need help What is the simplified form of the fifth root of x to the fourth power times the fifth root of x to the fourth power?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi i need help

OpenStudy (misty1212):

\[\sqrt[5]{x^4}\times \sqrt[5]{x^4}\]??

OpenStudy (anonymous):

yess

OpenStudy (misty1212):

first add up the exponents inside the radical (i.e. multiply) and get \[\sqrt[5]{x^8}\] but that is not the final answer

OpenStudy (anonymous):

10 /8 i add all the exponents

OpenStudy (misty1212):

since 5 goes in to 8 one time, with a remainder of 3, the final answer is \[x^1\sqrt[5]{x^3}\] or just \[x\sqrt[5]{x^3}\]

OpenStudy (misty1212):

you have it upside down dear, and the addition is not quite right \[\huge x^{\frac{4}{5}}\times x^{\frac{4}{5}}=x^{\frac{8}{5}}\] since \[\huge \frac{4}{5}+\frac{4}{5}=\frac{8}{5}=1\tfrac{3}{5}\]

OpenStudy (anonymous):

ohh i seee noww......

OpenStudy (anonymous):

Find the solution of 5 times the square root of the quantity of x plus 7 equals negative 10, and determine if it is an extraneous solution. x = −3; extraneous x = −3; not extraneous x = 11; extraneous x = 11; not extraneous

OpenStudy (misty1212):

that may be a good explanation of why the answer is \[x\sqrt[5]{x^3}\]

OpenStudy (anonymous):

my friend got A and i think is C who's right ?

OpenStudy (misty1212):

it is certainly extraneous, i.e. there is no solution so why should you have to try and find it i have no idea

OpenStudy (misty1212):

i guess they want you to square both sides \[25(x+7)=100\]

OpenStudy (misty1212):

divide by 25 get \[x+7=4\] so \[x=-3\] really it is a bad question though

OpenStudy (misty1212):

oh sorry, your friend is right

OpenStudy (anonymous):

oh i used the wrong operation :(

OpenStudy (anonymous):

Given the functions f(x) = −2x − 1 and g(x) = −3x + 4, which operation results in the smallest coefficient on the x term? Addition Subtraction Multiplication Two operations result in the same coefficient

OpenStudy (anonymous):

this is another one we are arguing about. she thinks its B i go with D

OpenStudy (misty1212):

ya'll both wrong

OpenStudy (anonymous):

ohhhhh why?

OpenStudy (misty1212):

add and get \(-5x+3\) subtract and get who knows because we don't know which one gets subtracted from which, but either way you will not get \(-5x\) you will get either \(-x\) or \(x\)

OpenStudy (misty1212):

oh wait, i didn't multiply yet

OpenStudy (anonymous):

i read to her what you wrote she thinks its A

OpenStudy (misty1212):

if you multiply you get \[6 x^2-5 x-4\]

OpenStudy (misty1212):

so the answer is actually D, two operations result in the same coefficient, addition and multiplication

OpenStudy (anonymous):

lol soo i was right

OpenStudy (misty1212):

let me guess FLVS?

OpenStudy (anonymous):

lol yeah im helping her. shes been struggling

OpenStudy (misty1212):

i can always tell flvs questions by how bad they are

OpenStudy (anonymous):

lol finally someone agrees... lol but thanks again for helping me i'll fan you okok